OpenStudy (anonymous):

Make two fractions with all the digits from 0 to 9. Use each digit once. They add up to exactly 1.

zepdrix (zepdrix):

Ooo that's an interesting problem c:

OpenStudy (anonymous):

Lol hard too . ^^ ;D

hartnn (hartnn):

not actually hard, just make sure both fractions simplify to 1/2

zepdrix (zepdrix):

I'm just like.. trying different combinations and seeing what works :p hmm.. how bout something over 908. 454/908. Hmmm no I used the 4 twice... How boutttttt something over 906. 453/906 = 1/2 Hmmmmmm we might be on the right track here :U maybeeeeee

zepdrix (zepdrix):

That would leave us with 1, 2, 7, 8 and we need to form another 1/2 with those numbers? Hmmmm

OpenStudy (anonymous):

Still confused. ;c

zepdrix (zepdrix):

Darn that didn't work :c hmmm...

zepdrix (zepdrix):

Too confusing? :o

OpenStudy (jack1):

1/2 + ...

OpenStudy (anonymous):

lol yahh tooo confusing ;/

OpenStudy (jack1):

3845 / 2(3845) = 3845 / 7690

OpenStudy (jack1):

done ;D

hartnn (hartnn):

and 3548/7096 will work too

OpenStudy (jack1):

true, nice method @hartnn ( with the : make them equal half)

hartnn (hartnn):

here's the whole list: 1/2+3485/6970, 1/2+3548/7096, 1/2+3845/7690, 1/2+4538/9076, 1/2+4685/9370, 1/2+4835/9670, 1/2+4853/9706, 1/2+4865/9730, 1/4+7365/9820, 1/6+7835/9402, 1/7+4362/5089, 2/4+3079/6158, 2/6+3190/4785, 2/7+5940/8316, 2/7+6810/9534, 2/9+5803/7461, 3/6+1485/2970, 3/6+2079/4158, 3/6+2709/5418, 3/6+2907/5814, 3/6+4851/9702, 3/127+496/508, 4/5+1278/6390, 4/5+1872/9360, 4/356+792/801, 5/104+693/728, 6/324+795/810, 6/534+792/801, 7/9+1208/5436, 7/9+1352/6084, 7/54+893/1026, 8/10+729/3645, 8/10+927/4635, 8/512+693/704, 9/12+876/3504, 9/351+684/702, 10/28+369/574, 10/45+287/369, 10/45+728/936, 10/96+473/528, 12/54+609/783, 12/60+748/935, 12/96+357/408, 12/96+735/840, 13/26+485/970, 13/52+678/904, 15/30+486/972, 16/32+485/970, 17/89+504/623, 18/90+276/345, 18/90+372/465, 19/57+308/462, 19/58+273/406, 21/96+375/480, 24/63+507/819, 24/96+531/708, 27/54+309/618, 27/81+306/459, 27/81+630/945, 29/58+307/614, 29/87+310/465, 31/62+485/970, 32/48+169/507, 32/80+417/695, 34/51+269/807, 34/578+96/102, 35/70+148/296, 35/70+481/962, 36/81+405/729, 36/81+540/972, 38/61+207/549, 38/76+145/290, 38/76+451/902, 38/95+426/710, 39/51+204/867, 39/65+284/710, 42/87+315/609, 45/61+208/793, 45/90+138/276, 45/90+186/372, 45/90+381/762, 46/92+185/370, 48/96+135/270, 48/96+351/702, 54/87+231/609, 56/84+109/327, 56/84+307/921, 56/428+93/107, 56/832+97/104, 57/92+140/368, 57/204+98/136, 59/236+78/104, 60/1245+79/83, 63/728+95/104, 70/96+143/528, 74/89+105/623, 87/435+96/120.

OpenStudy (jack1):

...what did u excel them?

zepdrix (zepdrix):

lol how did you generate that? XD

OpenStudy (anonymous):

WHOAAA THATS ALOT OF NUMBERS.. Lmao

hartnn (hartnn):

we can generate that by programming...but i googled :P

OpenStudy (jack1):

what's the formula dude if u dont mind me asking?

hartnn (hartnn):

lol, there's no formula here, just logic in coding.

OpenStudy (anonymous):

Thanks ^_^

hartnn (hartnn):

welcome ^_^